# BeNeRF: Neural Radiance Fields from a Single Blurry Image and Event Stream

### v1

@@ Abstract
Neural implicit representation of visual scenes has attracted a lot of attention in recent research of computer vision and graphics. Most prior methods focus on how to reconstruct 3D scene representation from a set of images. In this work, we demonstrate the possibility to recover the neural radiance fields (NeRF) from a single blurry image and its corresponding event stream. We model the camera motion with a cubic B-Spline in SE(3) space. Both the blurry image and the brightness change within a time interval, can then be synthesized from the 3D scene representation given the 6-DoF poses interpolated from the cubic B-Spline. Our method can jointly learn both the implicit neural scene representation and recover the camera motion by minimizing the differences between the synthesized data and the real measurements without pre-computed camera poses from COLMAP. We evaluate the proposed method with both synthetic and real datasets. The experimental results demonstrate that we are able to render view-consistent latent sharp images from the learned NeRF and bring a blurry image alive in high quality. Code and data are available at https://github.com/WU-CVGL/BeNeRF.
Figure 1: Given a single motion blurred image and its corresponding event stream, BeNeRF jointly learns neural radiance fields and camera motion trajectories. It synthesizes high-quality novel images along the camera trajectory, recovering a sharp and coherent video from the single blurry image.
@@ 2 Related Work
Neural Implicit Scene Representation. NeRF has attracted lots of attention due to its powerful ability of implicit 3D scene representation and novel view synthesis [33]. Many following works have been proposed to improve NeRF’s performance or extend NeRF to other fields. [35, 65, 8, 11] are proposed to speedup the training or rendering efficiency. [62, 29, 2] jointly trained NeRF with inaccurate camera poses. [31, 66, 6, 32, 37, 26] improved the performance of NeRF with degraded images, including noisy or few images etc. Some works are proposed to edit [60, 24, 67] or generate [41, 28] 3D scenes. There are also many approaches trying to extend NeRF to dynamic scenes [39, 42, 10, 40] or large scale scenes [56, 58]. Recently, several event based NeRF [49, 20, 44, 13] have also been proposed.
@@ 3.2 Camera Motion Trajectory Modeling
We use a differentiable cubic B-Spline in SE(3) space to better model the camera motion trajectory. The spline is represented by a set of learnable control knots $\mathbf{T}_{c_{i}}^{w}\in\mathbf{SE}(3)$ for $i=0,1,...,n$. $\mathbf{T}_{c_{i}}^{w}$ represents the $i^{th}$ control knot, which is defined as a transformation matrix from the camera coordinate frame to world frame. For brevity, we denote $\mathbf{T}_{c_{i}}^{w}$ with $\mathbf{T}_{i}$ for subsequent derivations. We assume the control knots are sampled with a uniform time interval $\Delta t$ and the trajectory starts from $t_{0}$. Spline with a smaller $\Delta t$ can represent a smoother motion, with an expense of more control knots to optimize. Since four consecutive control knots determine the value of the spline curve at a particular timestamp, we can thus compute the starting index of the four control knots for time $t$ by:
@@
where $\lfloor$*$\rfloor$ ∗ ⌋ is the floor operator. Then we can obtain the four control knots responsible for time $t$ as $\mathbf{T}_{k}$, $\mathbf{T}_{k+1}$, $\mathbf{T}_{k+2}$ and $\mathbf{T}_{k+3}$. We can further define $u=\frac{t-t_{0}}{\Delta t}-k$, where $u\in[0,1)$ to transform $t$ into a uniform time representation. Using this time representation and based on the matrix representation for the De Boor-Cox formula [45], we can write the matrix representation of a cumulative basis $\tilde{\mathbf{B}}(u)$ as
@@
$\mathbf{T}(u)=\mathbf{T}_{k}\cdot\prod_{j=0}^{2}\mathrm{exp}(\tilde{\mathbf{B}}(u)_{j+1}\cdot\boldsymbol{\Omega}_{k+j}),$ (5)
where $\tilde{\mathbf{B}}(u)_{j+1}$ denotes the $(j+1)^{th}$ element of the vector $\tilde{\mathbf{B}}(u)$, $\boldsymbol{\Omega}_{k+j}=\mathrm{log}(\mathbf{T}_{k+j}^{-1}\cdot\mathbf{T}_{k+j+1})$.
@@ 4.1 Experimental Setup
Table 2: Ablation studies on event accumulation time lengths. The experimental results demonstrate that the image quality gradually saturates as $\alpha$ increases to 0.25. For all experiments, we select $\alpha=0.1$ for event data accumulation.
@@
Table 3: Ablation studies on trajectory representations. The results demonstrate that cubic B-spline can better represent complex motions than linear interpolation, and deliver better performance in terms of the quality of the rendered image.
@@ 4.2 Ablation Study
Effect of time lengths for event accumulation. We study the effect of different time lengths $\Delta t$ for event accumulation as in Eq. (7). The timestamps of the event stream are normalized to a range of $(0,1)$ by its total time length. We choose different values from 0.05 to 0.25 for the ablation study. The experimental results are presented in Table 3, showing that the performance gradually saturates as $\alpha$ increases until to 0.25. In all experiments, we select $\alpha=0.1$.
@@ 4.3 Quantitative evaluations
We also compare against prior event-enhanced methods, such as EDI[38] and eSLNet[59]. The results in Table 5 demonstrates that our method has superior performance when compared to event-enhanced single image deblurring methods. eSLNet demonstrates poor generalization performance, since we are unable to fine-tune it on our evaluation datasets. It demonstrate the benefit on incorporating event streams to enhance single image deblurring task under the framework of NeRF.
@@ 4.4 Qualitative evaluations
The qualitative evaluation results are shown in Fig. 3 and Fig. 4 for both synthetic and real images respectively. The experimental results demonstrate that our method deliver better performance than prior methods . Even the image is severely blurred, our method still recovers it correctly. In particular, Fig. 3 shows that prior learning-based methods struggle to generalize to domain-shifted images. Notable the qualitative performance of EDI [38] deliver good qualitative performance on the synthetic datasets, which can be attributed to the perfect quality of the event data. Fig. 4 shows that our method outperforms all prior methods (even trained with multi-view images) on real noisy dataset, which clearly demonstrates the advantage of our method and the necessity to jointly optimize the camera motion and the implicit 3D representation.
@@ 5 Conclusion
In conclusion, we present a novel method to jointly recover the underlying 3D scene representation and camera motion trajectory from a single blurry image and its corresponding event stream. The method is formulated under the framework of neural radiance fields. Extensive experimental evaluations with both synthetic and real datasets demonstrate the superior performance of our method over prior works, even for those requiring multi-view images and longer event streams.
@@
Figure 4: Qualitative results of different methods on the real datasets. The experimental results demonstrate that our method delivers superior performance on the real DAVIS datasets from E2NeRF. We are even achieve better performance than prior methods requiring multi-view images and longer event data.

### v3

# https://github.com/wu-cvgl/BeNeRF

@@ Abstract
Implicit scene representation has attracted a lot of attention in recent research of computer vision and graphics. Most prior methods focus on how to reconstruct 3D scene representation from a set of images. In this work, we demonstrate the possibility to recover the neural radiance fields (NeRF) from a single blurry image and its corresponding event stream. To eliminate motion blur, we introduce event stream to regularize the learning process of NeRF by accumulating it into an image. We model the camera motion with a cubic B-Spline in SE(3) space. Both the blurry image and the brightness change within a time interval, can then be synthesized from the NeRF given the 6-DoF poses interpolated from the cubic B-Spline. Our method can jointly learn both the implicit scene representation and the camera motion by minimizing the differences between the synthesized data and the real measurements without any prior knowledge of camera poses. We evaluate the proposed method with both synthetic and real datasets. The experimental results demonstrate that we are able to render view-consistent latent sharp images from the learned NeRF and bring a blurry image alive in high quality.
Keywords:
Neural Radiance Fields Event Stream Pose Estimation Deblurring Novel View Synthesis 3D from a Single Image
11footnotetext: Equal contribution: {liwenpu, wangpeng}@westlake.edu.cn, pianwan@outlook.com44footnotetext: Corresponding author: Peidong Liu(liupeidong@westlake.edu.cn).
Figure 1: Given a single blurry image and its corresponding event stream, BeNeRF can synthesize high-quality novel images along the camera trajectory, recovering a sharp and coherent video from the single blurry image.
@@ 2 Related Work
Neural Implicit Scene Representation. NeRF has attracted lots of attention due to its powerful ability of implicit 3D scene representation and novel view synthesis [28]. Many following works have been proposed to improve NeRF’s performance or extend NeRF to other fields.[49, 24, 2] jointly trained NeRF with inaccurate camera poses. [26, 53, 6, 27, 32, 22] improved the performance of NeRF with degraded images, including noisy or few images etc. Recently, several event based NeRF [39, 17, 34, 10] have also been proposed.
@@ 3.2 Camera Motion Trajectory Modeling
We use a differentiable cubic B-Spline in SE(3) space to better model the camera motion trajectory. The spline is represented by a set of learnable control knots $\boldsymbol{T}_{c_{i}}^{w}\in\mathbb{SE}(3)$ for $i=0,1,...,n$. $\boldsymbol{T}_{c_{i}}^{w}$ represents the $i^{th}$ control knot, which is defined as a transformation matrix from the camera coordinate frame to world frame. For brevity, we denote $\boldsymbol{T}_{c_{i}}^{w}$ with $\boldsymbol{T}_{i}$ for subsequent derivations. We assume the control knots are sampled with a uniform time interval $\Delta t$ and the trajectory starts from $t_{0}$. Spline with a smaller $\Delta t$ can represent a smoother motion, with an expense of more control knots to optimize. Since four consecutive control knots determine the value of the spline curve at a particular timestamp, we can thus compute the starting index of the four control knots for time $t$ by:
@@
where $\lfloor$*$\rfloor$ ∗ ⌋ is the floor operator. Then we can obtain the four control knots responsible for time $t$ as $\boldsymbol{T}_{k}$, $\boldsymbol{T}_{k+1}$, $\boldsymbol{T}_{k+2}$ and $\boldsymbol{T}_{k+3}$. We can further define $u=\frac{t-t_{0}}{\Delta t}-k$, where $u\in[0,1)$ to transform $t$ into a uniform time representation. Using this time representation and based on the matrix representation for the De Boor-Cox formula [35], we can write the matrix representation of a cumulative basis $\mathcal{B}$(u) as
@@
$\boldsymbol{T}(u)=\boldsymbol{T}_{k}\cdot\prod_{j=0}^{2}\mathrm{exp}(\mathcal{B}(u)_{j+1}\cdot\boldsymbol{\Omega}_{k+j}),$ (5)
where $\mathcal{B}(u)_{j+1}$ denotes the $(j+1)^{th}$ element of the vector $\mathcal{B}(u)$, $\boldsymbol{\Omega}_{k+j}=\mathrm{log}(\boldsymbol{T}_{k+j}^{-1}\cdot\boldsymbol{T}_{k+j+1}))$.
@@ 4.1 Experimental Setup
Table 2: Ablation studies on event accumulation time lengths. The experimental results demonstrate that the image quality can be affected by the time length. For all experiments, we select $\alpha=0.1$ for event data accumulation.
@@
Table 3: Ablation studies on trajectory representations. The results demonstrate that cubic B-spline can deliver better performance than linear interpolation.
@@ 4.2 Ablation Study
Effect of time lengths for event accumulation. We study the effect of different time lengths $\Delta t$ for event accumulation as in Eq. (7). The timestamps of the event stream are normalized to a range of $(0,1)$ by its total time length. We choose different values from 0.05 to 0.25 for the ablation study. The experimental results are presented in Table 2, showing that the performance on Tanabata dataset gradually saturates as $\alpha$ increases until to 0.25, whereas on Livingroom dataset, performance initially improves with increasing $\alpha$ but subsequently declines. This may depend on the noise level in event stream of different time lengths. In all experiments, we select $\alpha=0.1$.
@@ 4.3 Quantitative evaluations
We also compare against prior event-enhanced single image deblurring methods, such as EDI[33] and eSLNet[47]. The results in Table 11 demonstrates that our method has superior performance when compared to them. eSLNet demonstrates poor generalization performance, since we are unable to fine-tune it on our evaluation datasets. It demonstrate the benefit on incorporating event streams to enhance single image deblurring task under the framework of NeRF.
@@ 4.4 Qualitative evaluations
The qualitative evaluation results are shown in Fig. 3 and Fig. 4 for both synthetic and real datasets respectively. The experimental results demonstrate that our method deliver better performance than prior methods even when the image is severely blurred. In particular, Fig. 3 shows that prior learning-based methods struggle to generalize to domain-shifted images. Notably, EDI[33] performs well on synthetic datasets due to the high quality of event data. Fig. 4 shows that our method outperforms all prior methods (even trained with multi-view images) on real noisy dataset, which demonstrates the advantage of our method and the necessity to jointly optimize the camera motion and the implicit 3D representation.
@@ 5 Conclusion
In conclusion, we present a novel method to jointly recover the underlying 3D scene representation and camera motion trajectory from a single blurry image and its corresponding event stream. Extensive experimental evaluations with both synthetic and real datasets demonstrate the superior performance of our method over prior works, even for those requiring multi-view images and longer event streams.
Acknowledgements. This work was supported in part by NSFC under Grant 62202389, in part by a grant from the Westlake University-Muyuan Joint Research Institute, and in part by the Westlake Education Foundation.
@@
Figure 4: Qualitative results of different methods on the real datasets. The experimental results demonstrate that our method delivers superior performance on the real DAVIS datasets from E2NeRF. We are even achieve better performance than prior methods requiring multi-view images and longer event stream.
Supplementary Material for “BeNeRF: Neural Radiance Fields from a Single Blurry Image and Event Stream”
[…]
Appendix 0.A Introduction
In this supplementary materials, we provide additional details and analysis on the ablation experiments. Furthermore, we present a more detailed quantitative comparison on synthetic datasets and an extensive qualitative comparison on both synthetic and real-world datasets by showcasing the deblurred images produced by various methods, facilitating a comprehensive evaluation. Finally, we provide the latent sharp videos recovered from a single blurred image using our method.
Appendix 0.B More ablation studies
Effect of event stream. We evaluate the effect of event stream on Livingroom dataset. Quantitative results are shown in Table 8 and qualitative results are presented in Fig. 5. It demonstrate that event streams can constrain the ill-posed problem caused by motion blur and only single view image as training data. Our method effectively incorporates event stream to guide NeRF to learn the correct underlying scene representation., significantly improving performance. Thus, the introduction of event streams is highly motivated.
Table 8: Ablation studies on event stream. The results demonstrate that leveraging event stream can dramatically boost the performance of BeNeRF. We validate that introducing event streams is an effective method to constrain the ill-posed problem caused by motion blur and limited to a single-view image as training data.
Livingroom
[…]
w/o event stream 24.40 .6612 .4712
w/ event stream 37.11 .9370 .0632
Figure 5: Qualitative results of ablation studies on event stream. Here we visualize how our method benefits from event stream. The qualitative results indicate that using event stream can lead to improved model performance.
Effect of the coarse-to-fine strategy for training. We experiment with the effect of introducing the coarse-to-fine strategy used in BARF[24] to train our model. The results in Table 9 show that incorporating the coarse-to-fine strategy slightly decreases model performance. It might be due that the coarse-to-fine strategy is more suit-able for multi-view case.
Table 9: Ablation studies on coarse-to-fine strategy for training. The results demonstrate that using coarse-to-fine strategy to train model does not further improve the performance
Livingroom
[…]
w/o coarse-to-fine 37.11 .9370 .0632
w/ coarse-to-fine 34.20 .8931 .1450
Appendix 0.C More quantitative results
Due to space constraints, we did not report the comparative results of SSIM metrics for each method on the synthetic dataset in the main text. In the supplementary material, we provide comprehensive quantitative comparison results, including metrics such as PSNR, SSIM, and LPIPS. The detailed quantitative results on our synthetic datasets are shown in Table 10, Table 11, Table 12. The experimental results presented in Table 10 demonstrate a significant superiority of our method over existing state-of-the-art single image deblurring methods. Similarly, the results in Table 11 exhibit superior performance of our method compared to event-enhanced single image deblurring methods. Additionally, the experimental results in Table 12 showcase the remarkable efficacy of our approach. Despite utilizing only a single blurred image and an event stream of a limited time interval, our method achieves performance comparable to E2NeRF [34], which employs multi-view images and a longer event stream, particularly in terms of the PSNR metric. Furthermore, our method even surpasses E2NeRF [34] in terms of the LPIPS metric.
Table 10: Detailed quantitative comparisons on single image deblurring with synthetic datasets. The results demonstrate that our method significantly performs better than prior learning-based methods in terms of image quality. For HINet and NAFNet, we tests pre-trained weights from both GoPro and REDS datasets(*).
PSNR $\uparrow$
Livingroom Whiteroom Pinkcastle Tanabata Outdoorpool Average
DeblurGANv2 [19] $29.26$ $27.64$ $23.16$ $20.09$ $26.89$ $25.41$
SRN-deblur [46] $30.86$ $27.59$ $23.12$ $19.89$ $27.79$ $25.85$
MPRNet [55] $28.57$ $26.49$ $21.60$ $18.20$ $27.02$ $24.38$
HINet [5] $28.56$ $26.27$ $21.91$ $18.59$ $26.70$ $24.41$
HINet* [5] $27.55$ $22.89$ $20.25$ $18.15$ $27.14$ $23.20$
NAFNet [4] $29.92$ $28.16$ $22.41$ $18.96$ $26.75$ $25.24$
NAFNet* [4] $28.18$ $23.67$ $20.85$ $18.38$ $27.52$ $23.72$
Restormer [54] $29.48$ $27.39$ $22.22$ $18.82$ $27.35$ $25.05$
BeNeRF $37.11$ $32.95$ $29.68$ $32.14$ $36.38$ $33.65$
SSIM $\uparrow$
Livingroom Whiteroom Pinkcastle Tanabata Outdoorpool Average
DeblurGANv2 [19] .$8121$$7235$$7043$$4964$$6123$6697.6697.6697.6697
SRN-deblur [46] .$8437$$7396$$7043$$5111$$6572$6912.6912.6912.6912
MPRNet [55] .$7937$$7301$$6547$$4258$$6253$6459.6459.6459.6459
HINet [5] .$7920$$6950$$6625$$4411$$6235$6428.6428.6428.6428
HINet* [5] .$7822$$6122$$6019$$4155$$6211$6066.6066.6066.6066
NAFNet [4] .$8306$$7874$$6896$$4665$$6255$6799.6799.6799.6799
NAFNet* [4] .$7991$$6422$$6175$$4230$$6407$6245.6245.6245.6245
Restormer [54] .$8262$$7314$$6803$$4596$$6352$6665.6665.6665.6665
BeNeRF .$9370$$8651$$8593$$9015$$9039$8934.8934.8934.8934
LPIPS $\downarrow$
Livingroom Whiteroom Pinkcastle Tanabata Outdoorpool Average
DeblurGANv2 [19] .2087 .1989 .2608 .3934 .3100 .2744
SRN-deblur [46] .2529 .2503 .3245 .4260 .3594 .3226
MPRNet [55] .2621 .2564 .3586 .4173 .3679 .3325
HINet [5] .2468 .2620 .3500 .4024 .3355 .3193
HINet* [5] .3327 .3602 .3789 .5265 .4397 .4076
NAFNet [4] .2268 .1991 .3058 .3908 .3280 .2901
NAFNet* [4] .3182 .3566 .3943 .5271 .4257 .4044
Restormer [54] .2391 .2493 .3373 .4248 .3664 .3234
BeNeRF .0632 .0788 .0761 .0515 .0677 .0675
Table 11: Detailed quantitative comparisons on event-enhanced single image deblurring with synthetic datasets. The results demonstrate that our method performs better than both EDI and eSLNet.
PSNR $\uparrow$
Livingroom Whiteroom Pinkcastle Tanabata Outdoorpool Average
eSLNet[47] 14.22 10.81 10.49 8.86 11.80 11.24
EDI[33] 32.61 30.33 27.24 24.87 31.64 29.34
BeNeRF 37.11 32.95 29.68 32.14 36.38 33.65
SSIM $\uparrow$
Livingroom Whiteroom Pinkcastle Tanabata Outdoorpool Average
eSLNet[47] .3527 .2156 .2903 .1658 .2181 .2485
EDI[33] .8871 .8152 .8356 .7564 .8044 .8197
BeNeRF .9370 .8651 .8593 .9015 .9039 .8934
LPIPS $\downarrow$
Livingroom Whiteroom Pinkcastle Tanabata Outdoorpool Average
eSLNet[47] .3981 .4236 .4902 .5067 .4676 .4572
EDI[33] .0904 .1020 .0779 .1039 .1409 .1030
BeNeRF .0632 .0788 .0761 .0515 .0677 .0675
Table 12: Detailed quantitative comparisons on NeRF-based image deblurring with synthetic datasets from E2NeRF. The results indicate that our method outperforms both NeRF and Deblur-NeRF, and exhibits performance comparable to E2NeRF in terms of the PSNR metric. Moreover, our method even surpasses E2NeRF with the LPIPS metric.
PSNR $\uparrow$
Chair Ficus Hotdog Lego Materials Mic Average
NeRF[28] 24.29 22.98 27.75 21.95 19.99 20.50 22.91
Deblur-NeRF[26] 25.87 22.86 24.62 24.47 20.54 11.92 21.71
E2NeRF[34] 31.28 30.00 34.34 28.11 27.27 27.60 29.77
BeNeRF 31.17 30.81 34.31 28.09 27.44 26.13 29.66
SSIM $\uparrow$
Chair Ficus Hotdog Lego Materials Mic Average
NeRF[28] .9357 .9023 .9546 .8548 .9108 .8854 .9072
Deblur-NeRF[26] .9373 .8982 .9396 .8756 .9012 .7249 .8795
E2NeRF[34] .9749 .9663 .9784 .9339 .9570 .9496 .9600
BeNeRF .9488 .9465 .9497 .8930 .9144 .9115 .9273
LPIPS $\downarrow$
Chair Ficus Hotdog Lego Materials Mic Average
NeRF[28] .1254 .1037 .1158 .2103 .1512 .1579 .1441
Deblur-NeRF[26] .2185 .1541 .2138 .2053 .2562 .3706 .2364
E2NeRF[34] .0608 .0362 .0660 .1078 .0919 .0724 .0725
BeNeRF .0500 .0299 .0539 .0745 .0708 .0738 .0588
Appendix 0.D More qualitative results
We conducted a detailed comparison of our proposed method with single-image deblurring methods and event-enhanced single-image deblurring methods on our proposed synthetic dataset, consisting of six scenes(i.e. Livingroom, Whiteroom, Pinkcastle, Tanabata and Outdoorpool). The motion blur images are synthesized by importing real camera motion trajectories from ETH3D [41]. Thorough qualitative comparison in Fig. 6 and Fig. 7 illustrate that even under severe motion blur conditions, our method can effectively recover sharp images, demonstrating significant superiority over existing state-of-the-art single-image deblurring methods. Since the event stream synthesized using ESIM [37] is monochannel, the images recovered using EDI[33] are grayscale.
To facilitate a comparative analysis with a NeRF-based image deblurring method leveraging multi-view information, we conducted a detailed evaluation on the synthetic dataset proposed by E2NeRF[34], which comprises six distinct scenes(i.e. Chair, Ficus, Hotdog, Lego, Materials, and Mic). Extensive qualitative comparisons in Fig. 8 and Fig. 9 demonstrate the superiority of our method, despite utilizing only a single blurred image and a short event stream, over methods that utilize multi-view images and long event streams.
Fig. 10 and Fig. 11 illustrate comprehensive comparisons on real-world datasets. The results demonstrate the superior performance of our method on real datasets, attributed to its enhanced capability in modeling the physical process of motion-blurred imaging.
Input Refer to caption
DeblurGANv2 Refer to caption
SRN Refer to caption
NAFNet Refer to caption
Restormer Refer to caption
EDI Refer to caption
BeNeRF Refer to caption
GT Refer to caption
Figure 6: Qualitative results of different methods with synthetic datasets. Detailed qualitative comparison for “Livingroom”, "Outdoorpool" and "Pinkcastle" scene of synthetic dataset.
Input Refer to caption
DeblurGANv2 Refer to caption
SRN Refer to caption
NAFNet Refer to caption
Restormer Refer to caption
EDI Refer to caption
BeNeRF Refer to caption
GT Refer to caption
Figure 7: Qualitative results of different methods with synthetic datasets. Detailed qualitative comparison for “Tanabata” and "Whiteroom" scene of synthetic dataset.
Input Refer to caption
NeRF Refer to caption
E2NeRF Refer to caption
BeNeRF Refer to caption
GT Refer to caption
Figure 8: Qualitative results of different methods with synthetic datasets proposed by E2NeRF. Detailed qualitative comparison for “Chair”, "Ficus" and "Hotdog" scene of synthetic dataset from E2NeRF.
Input Refer to caption
NeRF Refer to caption
E2NeRF Refer to caption
BeNeRF Refer to caption
GT Refer to caption
Figure 9: Qualitative results of different methods with synthetic datasets proposed by E2NeRF. Detailed qualitative comparison for “Lego”, "Materials" and "Mic" scene of synthetic dataset from E2NeRF.
Input Refer to caption
SRN Refer to caption
EDI Refer to caption
E2NeRF Refer to caption
BeNeRF Refer to caption
Figure 10: Qualitative results of different methods with real-world datasets. Detailed qualitative comparison for “Camera” and "Lego" scene of real-world dataset.
Input Refer to caption
SRN Refer to caption
EDI Refer to caption
E2NeRF Refer to caption
BeNeRF Refer to caption
Figure 11: Qualitative results of different methods with real-world datasets. Detailed qualitative comparison for “Letter”, "Plant" and "Toys" scene of real-world dataset.
Appendix 0.E Supplementary videos
To showcase the effectiveness of our approach, we provide a supplementary video illustrating its capability to recover high-quality latent sharp video from a single blurry image and corresponding event stream, which encapsulates rich temporal information. These videos are available on the project page. Furthermore, our results highlight the superior performance of our method in comparison to previous state-of-the-art approaches.